### v6

# Utilizing Large Language Models in an iterative paradigm with domain feedback for molecule optimization

Abstract
Molecule optimization is a critical task in drug discovery to optimize desired properties of a given molecule through chemical modification. Despite Large Language Models (LLMs) holding the potential to efficiently simulate this task by using natural language to direct the optimization, straightforwardly utilizing them shows limited performance. In this work, we facilitate utilizing LLMs in an iterative paradigm by proposing a simple yet highly effective domain feedback provider, namely $\text{Re}^{3}$. In detail, $\text{Re}^{3}$ harnesses an external toolkit, RDKit, to handle the molecule hallucination, if the modified molecule is chemically invalid. Otherwise, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective, followed by a retrieved example, to guide the LLM to refine the modified molecule. We conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{3}$ shows significant improvements. Particularly, for 20 single-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 16.95% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 6.04% and 5.25%.
Utilizing Large Language Models in an iterative paradigm
with domain feedback for molecule optimization
Khiem Le, Nitesh V. Chawla Department of Computer Science and Engineering University of Notre Dame, IN, USA {kle3, nchawla}@nd.edu
1 Introduction
Drug discovery, aimed at identifying molecules that can safely modulate the disease Bateman (2022), is a costly process that takes on average more than a decade to reach approval and release Berdigaliyev and Aljofan (2020). Fortunately, recent advancements in AI for molecular learning have been revolutionizing this process by remarkably accelerating many of its molecule-related tasks. While tasks such as molecule property prediction, as well as synthesis and retrosynthesis have been extensively investigated Zhao et al. (2023); Shi et al. (2023); Liu et al. (2024b); Le et al. (2024), molecule optimization, a critical task, remains largely under-explored.
Figure 1: An illustration of Refinement with $\text{Re}^{3}$. The domain feedback is generated from $\text{Re}^{3}$ to guide the LLM to refine the modified molecule.
Molecule optimization is the task of modifying a given molecule with the objective of achieving desired properties. Importantly, the modified molecule needs to retain a certain level of similarity to the original one, known as the similarity constraint of the modified molecule Jorgensen (2009). This task is routine in pharmaceutical companies and conventionally carried out by chemists through manual attempts Hoffer et al. (2018); de Souza Neto et al. (2020), which is not scalable. Given this context, the latest groundbreaking in AI, Large Language Models (LLMs), hold tremendous potential to efficiently simulate and scale this task by using natural language to direct the optimization Zhang et al. (2024). Unlike the traditional supervised training approach Yang et al. (2023); Xia et al. (2024), utilizing LLMs with outstanding abilities alleviates the need for labeled molecules Liu et al. (2023a), as well as exhibits zero-shot and open-vocabulary generalization beyond a pre-defined set of objectives, making them highly promising for this task.
As pioneers, Zhang et al. (2024) straightforwardly explore the capability of LLMs for solving this task using natural language to direct the optimization and notice limited performance. Moving further, motivated by the iterative refinement strategy employed by chemists, Liu et al. (2024a) propose to utilize LLMs in an iterative paradigm with domain feedback. Specifically, a domain feedback provider, ReDF verifies whether the modified molecule meets the objective, if not, ReDF retrieves from a prior database a molecule that is similar to the modified molecule and meets the objective, serving as an example to guide the LLM to refine the modified molecule. Notably, the domain feedback provider plays a vital role in the iterative paradigm Yang et al. (2022); Madaan et al. (2024). Although bringing improved performance, ReDF does not consider the molecule hallucination where the modified molecule is chemically invalid Guo et al. (2023), preventing the retrieving step. Moreover, by solely depending on a retrieved example as domain feedback, ReDF unexpectedly harms the similarity constraint of the modified molecule.
In this work, we propose an improved domain feedback provider, namely $\text{Re}^{3}$, which is carefully designed to facilitate utilizing LLMs in an iterative paradigm more effectively. Different from the previous ReDF, first of all, $\text{Re}^{3}$ takes into account the molecule hallucination, thereby actively promoting the retrieving step. To be more precise, at each iteration, the modified molecule is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013). If the modified molecule is chemically invalid, the concrete error message from RDKit such as the existence of an unclosed ring Schoenmaker et al. (2023); Skinnider (2024), is extracted as reliable domain feedback to inform the LLM to correct the modified molecule, in analogy to debugging Chen et al. (2023); Zhong et al. (2024). Otherwise, $\text{Re}^{3}$ verifies whether the modified molecule meets the objective, if not, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective to explicitly guide the LLM to refine the modified molecule. Additionally, as in the previous ReDF, a retrieved example is appended to further enrich the domain feedback. By diminishing the dependence on a retrieved example, $\text{Re}^{3}$ excels at satisfying the similarity constraint of the modified molecule, as we demonstrate along with improved performance.
We first revisit the evaluation metric used in previous works Liu et al. (2023b, 2024a), Hit ratio as the percentage of successful modifications, and identify a significant flaw in that failures involving modified molecules that are chemically invalid are excluded in calculating Hit ratio, resulting in an artificially high value. After fixing, we conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{3}$ shows significant improvements over baselines and excels at satisfying the similarity constraint of the modified molecule. Particularly, for 20 single-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 16.95% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 6.04% and 5.25%.
2 Methodology
Here we delve into the details of utilizing LLMs in an iterative paradigm integrated with the proposed domain feedback provider, $\text{Re}^{3}$.
Require: A given molecule $m$, an LLM $\mathcal{M}$,
and an external toolkit $\mathcal{T}$.
$\widehat{m}=\mathcal{M}(p||m)$triangleright initialization
For iteration $\in$ 0, …, N
       $\text{validity},\text{fb}_{\text{validity}}=\mathcal{T}(\widehat{m},\text{validity})$
       If not validity then
             $\widehat{m}=\mathcal{M}(p||m||\widehat{m}||\text{fb}_{\text{validity}})$
      Else
             If $E(C_{p}(\widehat{m}))=1$ then
                   break
            Else
                   $m_{e}=\text{retrieve}(\mathcal{D},\widehat{m})$triangleright (2)
                   $\widehat{m}=\mathcal{M}(p||m||\widehat{m}||\text{fb}||m_{e})$
Return $\widehat{m}$
Algorithm 1 Refinement with $\text{Re}^{3}$
2.1 Preliminaries
A given molecule $m$, represented in the SMILES string Weininger (1988). Molecule optimization is the task of modifying a given molecule with the objective of achieving desired properties, $p$. With an LLM $\mathcal{M}$, the modified molecule $\widehat{m}$ is formulated:
$\widehat{m}=\mathcal{M}$ $(p||m)\text{, }$ (1)
s.t. $E(C_{p}(\widehat{m}))=1\text{, }$
where ||||| | indicates concatenation, $E(C_{p}(\widehat{m}))\in\{0,1\}$ indicates whether $\widehat{m}$ meets the objective with its desired properties $C_{p}(\widehat{m})$.
Table 1: Results on single- and multi-property objectives as increase/decrease desired properties with 2 thresholds.
Desired properties thresholds Llama-3.1-8B Llama-3.1-70B
Vanilla Self-Feedback ReDF (3) Re³DF (3) Vanilla Self-Feedback ReDF (3) Re³DF (3)
Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑
+ LogP l 46.50 67.00 43.00 54.00 58.00 61.50 80.50 82.50 72.50 80.50 68.50 69.50 78.50 78.50 92.50 93.50
s 41.50 65.50 42.00 50.00 56.50 65.50 76.50 80.00 34.00 80.00 61.50 68.50 73.00 74.00 91.50 94.00
- LogP l 29.50 74.50 28.00 61.50 56.00 72.00 77.50 84.50 76.00 84.00 67.50 71.50 85.50 85.50 98.00 98.50
s 18.00 77.50 19.50 65.00 37.00 69.50 77.00 86.50 45.00 86.50 54.00 75.00 75.50 76.50 91.00 94.50
+ TPSA l 32.50 68.50 32.50 51.50 54.00 61.00 76.00 80.00 79.50 82.50 65.50 65.50 79.00 79.50 88.50 88.50
s 33.50 66.00 41.50 58.50 44.00 54.50 72.00 77.50 75.00 80.50 67.50 68.00 76.00 76.00 92.50 92.50
- TPSA l 25.50 78.50 32.00 67.50 53.50 66.50 85.50 88.00 38.50 79.00 36.00 62.00 71.00 72.00 92.50 96.00
s 17.50 78.50 27.00 64.00 43.50 66.50 84.50 89.50 31.50 82.00 39.00 68.00 72.50 74.50 93.50 95.50
+ HBDs l 23.50 71.50 31.00 50.00 49.50 60.00 75.50 81.50 65.00 74.50 39.50 43.50 78.50 78.50 93.00 93.50
s 09.00 67.50 18.00 53.50 35.50 52.50 69.00 77.50 18.00 79.50 33.50 52.50 62.00 63.50 78.50 82.00
- HBDs l 18.37 75.51 30.61 71.43 48.98 73.47 79.59 87.76 28.57 69.39 30.61 59.18 46.94 57.14 83.67 87.76
s 06.12 69.39 10.20 81.63 28.57 57.14 75.51 83.67 02.04 77.55 00.00 59.18 38.78 51.02 79.59 83.67
+ HBAs l 33.00 68.50 36.00 54.50 52.50 63.00 73.50 81.50 70.00 75.50 58.00 61.50 75.50 75.50 87.50 88.50
s 14.00 65.50 32.50 54.50 40.00 53.50 69.00 77.00 15.50 74.00 51.00 63.00 59.50 60.50 81.00 84.00
- HBAs l 30.61 79.59 40.82 67.35 63.27 63.27 93.88 97.96 40.82 77.55 44.90 61.22 75.51 79.59 89.80 93.88
s 12.24 69.39 16.33 59.18 46.94 69.39 81.63 89.80 10.20 71.43 20.41 55.10 71.43 75.51 83.67 89.80
+ Druglikeness l 24.00 81.00 10.50 61.50 46.50 67.50 74.50 84.00 39.50 83.50 28.00 70.00 64.50 78.50 85.50 90.50
s 05.00 77.00 04.00 67.00 25.00 66.50 50.00 82.50 07.00 84.00 04.50 68.50 35.50 74.50 63.50 91.50
- Druglikeness l 41.00 69.00 41.50 62.50 59.00 67.00 87.00 90.50 59.50 83.50 55.00 62.50 77.00 77.00 90.50 92.00
s 17.00 71.50 32.50 60.50 41.00 59.00 74.00 84.50 10.50 88.50 33.00 72.50 70.00 74.00 87.00 95.00
Average l 30.45 73.36 32.59 60.18 54.13 65.52 80.35 85.82 56.99 78.99 49.35 62.64 73.19 76.17 90.15 92.26
s 17.39 70.78 24.35 61.38 39.80 61.40 72.91 82.85 24.87 80.40 36.44 65.03 63.42 70.00 84.18 90.25
+ LogP & + TPSA l & l 20.50 70.00 26.50 48.50 34.00 56.50 54.50 68.00 18.00 78.50 18.00 52.50 47.00 63.00 59.50 84.00
s & s 16.50 62.00 20.50 44.00 25.50 51.00 43.50 62.50 07.00 80.00 19.50 64.00 40.00 64.00 56.50 83.00
+ LogP & - TPSA l & l 13.50 66.50 16.00 56.50 21.00 57.50 27.50 72.50 43.50 80.50 41.50 59.00 54.50 75.50 56.50 89.50
s & s 09.00 66.50 07.50 55.50 14.00 57.00 19.00 76.00 16.00 81.00 23.50 63.50 30.00 73.00 29.00 91.50
- LogP & + TPSA l & l 17.50 69.50 21.00 59.00 35.50 68.50 58.00 79.00 82.00 87.00 65.50 66.50 83.50 83.50 91.00 93.50
s & s 10.50 62.50 12.00 58.50 23.50 60.50 50.00 72.00 49.00 85.00 59.50 69.00 73.50 77.00 82.50 93.50
- LogP & - TPSA l & l 07.50 75.00 13.00 62.50 18.00 74.50 21.00 80.00 04.50 81.50 03.00 67.50 09.50 74.00 13.50 85.00
s & s 04.50 75.00 12.50 69.00 08.50 64.00 14.50 80.00 05.00 80.00 07.00 60.50 06.00 55.50 08.00 84.50
+ LogP & + HBDs l & l 17.00 70.00 16.00 52.50 26.50 52.50 52.00 64.50 03.50 82.50 19.50 52.50 49.50 67.00 60.50 84.00
s & s 09.00 64.00 14.00 48.50 22.50 52.00 43.00 65.50 02.00 78.00 05.50 55.00 38.50 71.00 47.00 81.50
+ LogP & - HBDs l & l 16.33 73.47 16.33 51.02 22.45 57.14 36.73 77.55 28.57 55.10 16.33 32.65 36.73 48.98 51.02 81.63
s & s 12.24 67.35 08.16 63.27 06.12 63.27 14.29 81.63 12.24 79.59 06.12 53.06 06.12 63.27 18.37 95.92
- LogP & + HBDs l & l 12.50 69.50 16.50 55.00 30.00 69.00 54.50 75.50 73.00 83.00 52.00 54.00 74.50 78.00 89.50 93.00
s & s 02.50 69.00 08.50 58.50 16.50 65.50 43.50 76.50 32.50 83.00 48.00 61.50 71.00 75.50 72.00 85.00
- LogP & - HBDs l & l 00.00 67.35 02.04 55.10 04.08 59.18 12.24 81.63 10.20 71.43 10.20 53.06 26.53 65.31 24.49 79.59
s & s 10.20 67.35 10.20 69.39 08.16 71.43 10.20 71.43 00.00 75.51 00.00 61.22 02.04 51.02 04.08 87.76
+ LogP & + HBAs l & l 15.00 59.00 17.50 42.50 29.00 53.50 45.50 65.00 15.50 84.00 30.00 57.50 50.00 65.50 52.00 76.00
s & s 05.50 62.00 15.00 52.50 24.50 56.50 41.50 58.50 06.00 78.00 23.50 61.00 37.00 56.00 56.00 85.00
+ LogP & - HBAs l & l 22.45 81.63 26.53 57.14 34.69 71.43 26.53 81.63 42.86 71.43 34.69 48.98 51.02 65.31 46.94 83.67
s & s 00.00 71.43 02.04 71.43 08.16 59.18 10.20 81.63 10.20 77.55 20.41 46.94 30.61 83.67 18.37 83.67
- LogP & + HBAs l & l 13.00 77.00 14.50 54.50 26.00 65.00 52.00 75.50 57.50 85.00 60.50 68.00 80.50 84.00 78.00 90.50
s & s 03.00 66.00 07.00 59.00 12.50 62.00 41.50 70.00 24.00 88.00 50.00 67.00 61.00 70.50 69.50 89.50
- LogP & - HBAs l & l 08.16 77.55 14.29 57.14 22.45 67.35 18.37 79.59 04.08 79.59 06.12 51.02 10.20 57.14 08.16 81.63
s & s 10.20 75.51 22.45 77.55 14.29 63.27 12.24 87.76 02.04 75.51 02.04 59.18 10.20 65.31 06.12 79.59
+ LogP & + Druglikeness l & l 10.00 69.50 05.50 54.50 27.50 58.00 42.50 72.00 12.50 83.50 07.50 63.50 40.00 78.50 46.50 86.00
s & s 01.00 66.50 00.50 55.50 08.00 68.00 31.00 73.00 03.00 87.50 02.00 63.50 17.50 82.50 23.50 85.00
+ LogP & - Druglikeness l & l 38.00 68.00 34.00 55.00 44.00 68.50 49.50 76.50 72.00 85.00 54.50 57.50 72.00 81.00 77.50 88.50
s & s 16.00 77.50 22.00 62.00 23.00 69.50 20.00 83.50 14.50 87.00 37.00 63.00 17.50 72.50 25.00 88.00
- LogP & + Druglikeness l & l 18.00 79.00 13.50 60.00 36.00 73.50 51.00 76.50 31.50 84.50 09.00 65.00 46.50 75.00 62.00 85.00
s & s 02.50 76.00 01.00 65.50 15.50 66.00 36.50 77.00 02.50 83.50 02.50 70.50 21.00 79.50 33.50 84.00
- LogP & - Druglikeness l & l 19.50 77.50 22.00 56.00 21.50 74.00 19.50 76.00 42.00 72.50 48.00 55.50 50.00 68.00 61.50 90.00
s & s 32.00 79.00 34.00 75.50 38.00 77.00 34.50 82.50 05.50 83.00 23.00 58.50 12.50 70.00 09.00 87.00
Average l & l 15.56 71.91 17.20 54.81 27.04 64.13 38.84 75.09 33.83 79.07 29.77 56.54 48.87 70.61 54.91 85.72
s & s 09.04 69.23 12.33 61.60 16.80 62.88 29.09 74.97 11.97 81.39 20.60 61.09 29.65 69.39 34.90 86.53
2.2 Refinement with $\text{Re}^{3}$
After initialization, $\widehat{m}$ is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013), a widely-used cheminformatics toolkit. If $\widehat{m}$ is chemically invalid, the concrete error message from RDKit is extracted as reliable domain feedback to inform the LLM $\mathcal{M}$ to correct $\widehat{m}$. Otherwise, $\text{Re}^{3}$ verifies whether $\widehat{m}$ meets the objective by checking whether $E(C_{p}(\widehat{m}))=1$, if not, its desired properties $C_{p}(\widehat{m})$, are computed and compared to the original one, establishing reliable domain feedback, denoted fb, with correct direction and distance towards the objective, $|C_{p}(\widehat{m})-C_{p}(m)|$, to explicitly guide the LLM $\mathcal{M}$ to refine $\widehat{m}$. Additionally, $\text{Re}^{3}$ retrieves from a prior database $\mathcal{D}$ a molecule $m_{e}$ that is similar to $\widehat{m}$ and meets the objective, serving as an example to further enrich the domain feedback:
$m_{e}$ $=\text{retrieve}(\mathcal{D},\widehat{m})\text{, }$ (2)
$=\underset{m_{e}^{\prime}\in\mathcal{D}}{\text{argmax}}\left\langle m_{e}^{\prime},\widehat{m}\right\rangle\wedge E(C_{p}(m_{e}^{\prime}))\text{, }$
where $\left\langle m_{e}^{\prime},\widehat{m}\right\rangle$ is the Tanimoto similarity Bajusz et al. (2015) between a molecule $m_{e}^{\prime}$ in $\mathcal{D}$ and $\widehat{m}$. Subsequently, the described procedure is iterated for a pre-defined maximum number of iterations N or either can be early stopped if $\widehat{m}$ meets the objective, as precisely depicted in Algorithm 1.
3 Experiments
Table 2: The similarity constraint of the modified molecule on single- and multi-property objectives as increase/decrease desired properties with 2 thresholds, quantified by the Tanimoto similarity to the original one.
Desired properties thresholds Llama-3.1-8B Llama-3.1-70B
ReDF (3) Re³DF (3) ReDF (3) Re³DF (3)
similarity↑ similarity↑ similarity↑ similarity↑
+ LogP l 46.61 47.53 70.46 72.48
s 39.23 41.36 51.86 63.93
- LogP l 46.64 45.12 71.70 73.90
s 38.40 39.41 60.82 66.20
+ TPSA l 46.52 46.46 72.14 71.56
s 43.66 41.70 69.55 70.99
- TPSA l 39.37 39.02 53.29 54.94
s 33.95 32.22 45.62 47.76
+ HBDs l 42.42 40.14 68.81 72.72
s 30.49 32.07 37.60 55.13
- HBDs l 39.47 41.55 53.78 66.02
s 29.15 33.08 31.07 41.18
+ HBAs l 46.48 48.58 68.00 71.41
s 31.25 33.34 39.01 56.93
- HBAs l 37.99 42.74 53.44 58.59
s 29.87 35.88 38.18 42.44
+ Druglikeness l 42.69 42.26 52.43 59.07
s 39.76 35.05 44.70 53.80
- Druglikeness l 49.21 50.73 66.73 71.50
s 37.33 39.45 40.14 57.05
Average l 43.74 44.41 63.08 67.22
s 35.31 36.36 45.85 55.54
+ LogP & + TPSA l & l 36.44 34.08 41.95 58.08
s & s 33.45 34.08 33.34 52.34
+ LogP & - TPSA l & l 39.50 40.53 50.67 53.55
s & s 38.89 33.89 43.14 47.61
- LogP & + TPSA l & l 37.24 40.56 68.67 69.17
s & s 38.08 36.10 57.70 64.62
- LogP & - TPSA l & l 35.58 39.90 41.54 59.90
s & s 34.54 38.00 42.21 46.96
+ LogP & + HBDs l & l 31.96 32.26 34.14 54.52
s & s 30.93 27.87 33.34 46.84
+ LogP & - HBDs l & l 32.86 34.10 47.84 54.00
s & s 24.14 25.55 36.49 41.97
- LogP & + HBDs l & l 37.44 36.03 64.67 65.43
s & s 34.68 28.97 42.96 55.90
- LogP & - HBDs l & l 40.35 33.37 41.29 51.94
s & s 26.80 24.70 35.17 44.37
+ LogP & + HBAs l & l 35.54 35.36 40.56 59.09
s & s 32.32 29.02 33.40 49.13
+ LogP & - HBAs l & l 37.79 34.23 50.53 56.85
s & s 30.66 27.63 38.13 41.31
- LogP & + HBAs l & l 35.48 39.61 61.29 65.75
s & s 36.03 30.90 39.88 53.39
- LogP & - HBAs l & l 33.34 36.12 43.95 54.59
s & s 28.13 27.92 40.58 43.06
+ LogP & + Druglikeness l & l 39.01 40.55 39.44 56.45
s & s 37.16 32.55 45.43 57.88
+ LogP & - Druglikeness l & l 40.98 37.43 59.33 63.13
s & s 28.21 24.93 45.85 55.32
- LogP & + Druglikeness l & l 43.95 40.57 49.45 62.56
s & s 39.95 32.33 47.18 59.21
- LogP & - Druglikeness l & l 39.90 43.01 59.72 65.42
s & s 19.85 22.15 42.56 53.29
Average l & l 37.34 37.36 49.69 59.40
s & s 32.11 29.79 41.08 50.83
Here we conduct experiments to evaluate the effectiveness of the proposed domain feedback provider, $\text{Re}^{3}$ in integrating with LLMs in an iterative paradigm. Following previous works Liu et al. (2023b, 2024a), 200 molecules are sampled from the ZINC database Irwin et al. (2012) as given molecules. Using 5 high-level molecular properties such as LogP as desired properties, we adopt single- and multi-property objectives as increase/decrease (+/-) desired properties with 2 thresholds. Details of the set of examined objectives and prompts in natural language are listed in Appendix A1. Throughout experiments, we utilize the latest and most capable herd of open-sourced LLMs, Llama-3.1 Dubey et al. (2024) with 2 model sizes 8B and 70B parameters. Straightforwardly prompting LLMs is a vanilla baseline Zhang et al. (2024). Furthermore, for stronger baselines, we compare $\text{Re}^{3}$ in integrating with LLMs in a 3-iteration (3) iterative paradigm with Self-Feedback Madaan et al. (2024) in which LLMs themselves act as the domain feedback provider, where prompt for acquiring domain feedback is adapted from the primary paper, and especially compare with the previous ReDF. As in previous works Liu et al. (2023b, 2024a), Hit ratio which is the percentage of successful modifications, is used as the evaluation metric. However, we identify a significant flaw in that failures involving modified molecules that are chemically invalid are intentionally skipped and excluded in calculating Hit ratio, resulting in an artificially high value. In light of this, to reflect the actual value, we purely calculate Hit ratio as the percentage of successful modifications without excluding modified molecules that are chemically invalid and separately report the validity ratio as an additional evaluation metric Edwards et al. (2022) for more transparency.
Table 1 displays results across both single- and multi-property objectives with 2 thresholds. In general, utilizing LLMs in an iterative paradigm even with Self-Feedback Madaan et al. (2024) whose domain feedback lacks fidelity, shows better performance than the vanilla baseline. In a particular comparison with the previous ReDF, for 20 single-property objectives, $\text{Re}^{3}$ with Llama-3.1-70B yields best results on 19/20 objectives, enhancing Hit ratio by 16.95% and 20.76% while raising the validity ratio by 16.09% and 20.25% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ with Llama-3.1-70B yields best results on 18/32 objectives, enhancing Hit ratio by 6.04% and 5.25% while raising the validity ratio by 15.11% and 17.14%. Along with improved performance, Table 2 demonstrates a crucial advantage of $\text{Re}^{3}$ in preserving the similarity constraint of the modified molecule compared to the previous ReDF, quantified by the Tanimoto similarity to the original one, emphasizing the importance of reliable domain feedback. Particularly, for 20 single-property objectives, $\text{Re}^{3}$ with Llama-3.1-70B boosts the Tanimoto similarity by 4.14 and 9.68 under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ with Llama-3.1-70B boosts the Tanimoto similarity by 9.71 and 9.74.
4 Conclusion
In this work, we make focused contributions on utilizing LLMs with outstanding abilities for solving molecule optimization, a critical task in drug discovery, using natural language to direct the optimization. To this end, we propose an improved domain feedback provider, namely $\text{Re}^{3}$, which is carefully designed to facilitate utilizing LLMs in an iterative paradigm more effectively. For evaluation, after fixing a significant flaw in calculating Hit ratio, we conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{3}$ shows significant improvements over baselines and excels at satisfying the similarity constraint of the modified molecule. Besides the main experiments, we conduct an analysis to understand the impact of the maximum number of iterations N in the iterative paradigm and observe that $\text{Re}^{3}$ is beneficial from growing N and consistently outperforms the previous ReDF across both single- and multi-property objectives with 2 thresholds, as presented in Table 4, Appendix A2.
Limitations
First of all, although it does not require any training, we admit that the strategy of utilizing LLMs in an iterative paradigm requires prompting LLMs over multiple iterations, obviously leading to multiplying prompting costs Samsi et al. (2023). However, this limitation is acceptable and compensated by bringing improved performance. With the awareness that the domain feedback provider plays a vital role in the iterative paradigm, we propose an improved domain feedback provider, namely $\text{Re}^{3}$. By establishing reliable domain feedback with correct direction and distance towards the objective, followed by a retrieved example, to guide the LLM to refine the modified molecule, $\text{Re}^{3}$ outperforms the previous ReDF across both single- and multi-property objectives with 2 thresholds. Albeit containing correct direction and distance towards the objective, domain feedback from $\text{Re}^{3}$ currently lacks correct actions towards the objective which involve knowledge of molecular properties Fang et al. (2023); Hoang et al. (2024), are presumably useful but hard to establish.
Appendix A1 Details of Experiments
To conduct experiments, a sufficient number of 200 molecules are sampled from the ZINC database Irwin et al. (2012) as given molecules. Then, 5 high-level molecular properties, i.e., LogP, TPSA, Hydrogen bond donors (HBDs), Hydrogen bond acceptors (HBAs), and Druglikeness are selected as desired properties. As stated in previous works Liu et al. (2023b, 2024a), these high-level molecular properties are selected since they have an ambiguous nature and involve indeterministic answers, serving as a source of inspiration for chemists. Another criterion for selection is that they can be computed by computational software, ensuring evaluation at high volume. Using these molecular properties, we adopt single- and multi-property objectives as increase/decrease (+/-) desired properties with 2 thresholds, as expressed in Table 3. These objectives are then wrapped into natural language with the template of prompts kept naive as listed in Box 1. As discussed earlier, straightforwardly prompting LLMs is a vanilla baseline Zhang et al. (2024). Furthermore, for stronger baselines, we compare $\text{Re}^{3}$ in integrating with LLMs in a 3-iteration (3) iterative paradigm with Self-Feedback Madaan et al. (2024) in which LLMs themselves act as the domain feedback provider, where prompt for acquiring domain feedback is adapted from the primary paper and are listed in Box 2, and especially compare with the previous ReDF. Different from the previous ReDF, $\text{Re}^{3}$ takes into account the molecule hallucination. To be more precise, at each iteration, the modified molecule is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013). If the modified molecule is chemically invalid, the concrete error message from RDKit which belongs to one of six categories: syntax error, parentheses error (extra open parentheses or extra close parentheses), unclosed ring, aromaticity error (a combination of non-ring atom marked aromatic and kekulization errors), bond already exists, and valence error Schoenmaker et al. (2023), is extracted as reliable domain feedback to inform the LLM to correct the modified molecule. By establishing reliable domain feedback with correct direction and distance towards the objective, followed by a retrieved example, to guide the LLM to refine the modified molecule as illustrated in Listing 1, $\text{Re}^{3}$ outperforms the previous ReDF across both single- and multi-property objectives with 2 thresholds.
Table 3: Detailed information on the set of examined objectives with single- and multi-property objectives as increase/decrease desired properties with 2 thresholds.
Desired properties thresholds magnitudes of thresholds
+ LogP l 00.0
s 00.5
- LogP l 00.0
s 00.5
+ TPSA l 00.0
s 10.0
- TPSA l 00.0
s 10.0
+ HBDs l 00.0
s 01.0
- HBDs l 00.0
s 01.0
+ HBAs l 00.0
s 01.0
- HBAs l 00.0
s 01.0
+ Druglikeness l 00.0
s 00.1
- Druglikeness l 00.0
s 00.1
Average l -
s -
+ LogP & + TPSA l & l 00.0 & 00.0
s & s 00.5 & 10.0
+ LogP & - TPSA l & l 00.0 & 00.0
s & s 00.5 & 10.0
- LogP & + TPSA l & l 00.0 & 00.0
s & s 00.5 & 10.0
- LogP & - TPSA l & l 00.0 & 00.0
s & s 00.5 & 10.0
+ LogP & + HBDs l & l 00.0 & 00.0
s & s 00.5 & 01.0
+ LogP & - HBDs l & l 00.0 & 00.0
s & s 00.5 & 01.0
- LogP & + HBDs l & l 00.0 & 00.0
s & s 00.5 & 01.0
- LogP & - HBDs l & l 00.0 & 00.0
s & s 00.5 & 01.0
+ LogP & + HBAs l & l 00.0 & 00.0
s & s 00.5 & 01.0
+ LogP & - HBAs l & l 00.0 & 00.0
s & s 00.5 & 01.0
- LogP & + HBAs l & l 00.0 & 00.0
s & s 00.5 & 01.0
- LogP & - HBAs l & l 00.0 & 00.0
s & s 00.5 & 01.0
+ LogP & + Druglikeness l & l 00.0 & 00.0
s & s 00.5 & 00.1
+ LogP & - Druglikeness l & l 00.0 & 00.0
s & s 00.5 & 00.1
- LogP & + Druglikeness l & l 00.0 & 00.0
s & s 00.5 & 00.1
- LogP & - Druglikeness l & l 00.0 & 00.0
s & s 00.5 & 00.1
Average l & l -
s & s -
Table 4: Results on single- and multi-property objectives as increase/decrease desired properties with 2 thresholds.
Desired properties thresholds Llama-3.1-8B Llama-3.1-70B
ReDF (4) ReDF (5) Re³DF (4) Re³DF (5) ReDF (4) ReDF (5) Re³DF (4) Re³DF (5)
Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑ Hit↑ validity↑
+ LogP l 59.00 62.50 60.00 61.50 83.00 84.50 86.50 88.00 78.50 78.50 78.50 78.50 94.00 94.50 95.50 95.50
s 50.00 57.00 61.50 67.00 78.50 82.50 88.00 89.00 75.50 76.50 77.00 77.50 91.00 94.00 95.00 96.00
- LogP l 51.50 64.50 54.50 67.50 83.00 87.00 86.00 93.50 85.00 85.00 85.00 85.00 97.00 97.50 99.00 99.00
s 44.50 62.50 43.00 60.00 77.00 87.00 84.00 89.00 74.00 75.00 72.50 74.00 92.50 94.00 95.50 95.50
+ TPSA l 59.00 63.00 54.00 58.00 81.50 83.50 81.00 82.50 81.00 81.00 81.00 81.00 93.50 93.50 94.00 94.00
s 53.50 58.00 46.50 55.50 79.50 84.00 85.00 86.00 76.00 76.00 76.00 76.00 94.00 94.50 95.00 95.00
- TPSA l 58.00 65.00 66.00 72.50 87.00 90.00 88.00 89.00 71.50 72.50 69.50 70.50 94.00 95.00 96.50 97.50
s 46.50 58.50 51.50 64.50 85.50 91.00 88.00 89.50 73.00 75.50 74.50 77.00 94.00 96.50 95.00 97.00
+ HBDs l 54.00 61.50 51.50 58.00 82.50 84.50 82.50 85.00 78.50 78.50 78.50 78.50 93.00 93.50 92.50 93.00
s 46.50 63.00 43.00 55.00 68.50 77.50 70.50 80.50 62.50 63.00 62.50 63.00 86.00 88.00 85.00 88.50
- HBDs l 51.02 71.43 61.22 73.47 75.51 85.71 75.51 83.67 59.18 63.27 59.18 61.22 89.80 89.80 85.71 87.76
s 36.73 69.39 36.73 51.02 73.47 87.76 79.59 91.84 38.78 51.02 51.02 55.10 85.71 89.80 95.92 95.92
+ HBAs l 53.50 61.50 51.00 57.50 77.00 80.50 80.50 85.00 75.50 75.50 75.50 75.50 88.00 88.50 91.50 92.00
s 43.50 53.50 34.00 49.00 71.00 77.50 74.00 77.50 56.50 56.50 56.50 56.50 88.00 90.00 85.00 88.50
- HBAs l 63.27 63.27 63.27 63.27 81.63 89.80 89.80 91.84 77.55 77.55 77.55 77.55 91.84 91.84 91.84 95.92
s 30.61 61.22 42.86 53.06 87.76 91.84 83.67 87.76 67.35 73.47 77.55 79.59 89.80 95.92 93.88 95.92
+ Druglikeness l 51.00 68.00 51.50 66.00 77.50 85.50 82.00 88.50 66.00 79.50 68.00 76.50 92.50 95.50 89.00 92.00
s 31.50 72.00 29.00 67.50 61.00 87.50 58.50 83.00 36.00 73.00 41.00 73.00 66.00 92.00 65.50 92.00
- Druglikeness l 62.00 67.00 66.00 68.50 86.50 89.00 82.50 87.50 77.00 77.00 77.00 77.00 95.00 96.50 96.50 97.50
s 42.50 58.50 46.50 59.50 83.50 89.00 76.50 83.50 60.00 64.50 67.00 69.50 90.00 94.50 93.00 94.50
Average l 56.23 64.77 57.90 64.62 81.51 86.00 83.43 87.45 74.97 76.83 74.97 76.13 92.86 93.61 93.21 94.42
s 42.58 61.36 43.46 58.21 76.57 85.56 78.78 85.76 61.96 68.45 65.56 70.12 87.70 92.92 89.88 93.88
+ LogP & + TPSA l & l 34.00 56.50 34.00 56.50 57.00 66.00 53.00 65.00 44.00 63.00 44.00 63.00 58.00 78.50 58.00 79.50
s & s 25.50 51.00 25.50 51.00 48.00 61.50 51.50 63.00 40.00 64.00 40.00 64.00 67.50 87.00 54.50 81.50
+ LogP & - TPSA l & l 21.00 57.50 21.00 57.50 30.00 71.50 29.50 71.00 54.50 75.50 54.50 75.50 58.50 88.50 59.00 86.00
s & s 14.00 57.00 14.00 57.00 15.00 72.00 14.50 73.00 30.00 73.00 30.00 73.00 30.00 86.50 29.50 88.00
- LogP & + TPSA l & l 35.50 68.50 35.50 68.50 55.50 78.00 57.00 73.50 83.50 83.50 83.50 83.50 91.50 94.00 92.00 93.00
s & s 23.50 60.50 23.50 60.50 48.50 69.50 51.00 70.50 73.50 77.00 73.50 77.00 80.00 91.00 79.00 92.50
- LogP & - TPSA l & l 18.00 74.50 18.00 74.50 16.50 73.00 17.00 75.50 09.50 73.50 09.50 73.50 16.50 82.50 17.50 86.50
s & s 08.50 64.00 08.50 64.00 07.00 75.00 9.50 75.50 06.00 55.50 06.00 55.50 06.00 86.00 06.00 82.50
+ LogP & + HBDs l & l 26.50 52.50 26.50 52.50 52.00 68.00 50.50 68.50 49.00 67.00 49.00 67.00 46.00 76.50 53.50 82.00
s & s 22.50 52.00 22.50 52.00 43.00 62.00 38.00 62.50 35.50 71.00 35.50 71.00 48.50 83.50 46.50 79.50
+ LogP & - HBDs l & l 22.45 57.14 22.45 57.14 16.33 77.55 44.90 79.59 36.73 48.98 36.73 48.98 32.65 67.35 46.94 81.63
s & s 06.12 63.27 06.12 63.27 16.33 81.63 08.16 61.22 06.12 63.27 06.12 63.27 12.24 83.67 20.41 79.59
- LogP & + HBDs l & l 30.00 69.00 30.00 69.00 58.50 75.00 54.00 74.00 74.50 78.00 74.50 78.00 87.00 91.50 86.00 89.00
s & s 16.50 65.50 16.50 65.50 44.50 76.00 45.50 75.00 71.00 75.50 71.00 75.50 72.00 89.50 75.50 83.50
- LogP & - HBDs l & l 04.08 59.18 04.08 59.18 20.41 87.76 16.33 83.67 26.53 65.31 26.53 65.31 26.53 79.59 16.33 75.51
s & s 08.16 71.43 08.16 71.43 12.24 83.67 10.20 77.55 02.04 51.02 02.04 51.02 04.08 89.80 00.00 77.55
+ LogP & + HBAs l & l 29.00 53.50 29.00 53.50 51.00 67.50 50.00 65.00 49.00 65.00 49.00 65.00 58.00 81.50 59.00 84.00
s & s 24.50 56.50 24.50 56.50 44.00 64.00 46.50 62.50 38.00 57.00 38.00 57.00 56.50 81.00 59.50 75.50
+ LogP & - HBAs l & l 34.69 71.43 34.69 71.43 42.86 87.76 16.33 63.27 51.02 65.31 51.02 65.31 44.90 79.59 55.10 91.84
s & s 08.16 59.18 08.16 59.18 06.12 79.59 10.20 63.27 30.61 83.67 30.61 83.67 22.45 87.76 16.33 75.51
- LogP & + HBAs l & l 26.00 65.00 26.00 65.00 52.00 72.50 43.00 69.00 80.50 84.00 80.50 84.00 81.00 91.50 81.00 89.50
s & s 12.50 62.00 12.50 62.00 36.50 69.00 37.00 71.00 61.00 70.50 61.00 70.50 59.50 84.50 66.00 87.50
- LogP & - HBAs l & l 22.45 67.35 22.45 67.35 24.49 91.84 16.33 81.63 10.20 57.14 10.20 57.14 18.37 79.59 20.41 83.67
s & s 14.29 63.27 14.29 63.27 18.37 79.59 16.33 75.51 10.20 65.31 10.20 65.31 10.20 87.76 08.16 91.84
+ LogP & + Druglikeness l & l 27.50 58.00 27.50 58.00 40.00 71.00 42.50 68.50 40.00 78.50 40.00 78.50 47.50 83.00 45.50 88.00
s & s 08.00 68.00 08.00 68.00 24.00 68.50 33.00 70.50 17.50 82.50 17.50 82.50 26.50 81.00 27.00 83.00
+ LogP & - Druglikeness l & l 44.00 68.50 44.00 68.50 50.50 77.50 50.50 77.50 72.00 80.50 72.00 80.50 69.50 84.00 76.00 88.50
s & s 23.00 69.50 23.00 69.50 21.50 76.00 24.50 83.00 17.50 72.50 17.50 72.50 16.50 89.00 19.00 89.00
- LogP & + Druglikeness l & l 36.00 73.50 36.00 73.50 53.00 76.50 52.00 75.00 46.50 75.00 46.50 75.00 59.00 87.00 60.00 87.00
s & s 15.50 66.00 15.50 66.00 34.50 75.00 33.00 77.00 21.00 79.50 21.00 79.50 35.00 86.00 34.00 82.50
- LogP & - Druglikeness l & l 21.50 74.00 21.50 74.00 22.50 75.00 24.00 81.00 50.00 67.50 50.00 67.50 60.50 85.50 59.00 89.00
s & s 38.00 77.00 38.00 77.00 33.00 85.50 31.50 82.50 12.50 70.00 12.50 70.00 09.00 88.00 12.50 86.00
Average l & l 27.04 64.13 27.04 64.13 40.16 76.03 38.56 73.23 48.59 70.48 48.59 70.48 53.47 83.13 55.33 85.92
s & s 16.80 62.88 16.80 62.88 28.29 73.66 28.77 71.47 29.53 69.45 29.53 69.45 34.75 86.37 34.62 83.47
Box 1: Template of prompts for initialization. With a given molecule [$m$], modify it to [increase/decrease] its [desired properties] by at least [magnitudes of thresholds]. Importantly, the modified molecule should retain a high similarity to the original one. Respond with only the SMILES string of the modified molecule. No explanation is needed.
Box 2: Prompt for acquiring domain feedback in Self-Feedback Madaan et al. (2024). Assess the modified molecule on its desired properties according to the objective then provide domain feedback to refine the modified molecule. Start your domain feedback with the phrase "Desired properties: ".
Improve the modified molecule based on the domain feedback: Desired properties: The modified molecule has [desired properties] of [$C_{p}(\widehat{m})$], which [increased/decreased] by [$|C_{p}(\widehat{m})-C_{p}(m)|$] compared to the original one. Therefore, the modified molecule does not meet the objective. For your reference, we found a molecule [$m_{e}$] that is similar to the modified molecule and meets the objective. Respond with only the SMILES string of the modified molecule. No explanation is needed.
Listing 1: The domain feedback from $\text{Re}^{3}$ to guide the LLM to refine the modified molecule.
Hit $\text{Hit ratio}=\frac{\text{number of successful modifications}}{\text{numberof given molecules}}\%$ (3)
As mentioned, we identify a significant flaw in that failures involving modified molecules that are chemically invalid are intentionally skipped and excluded in calculating Hit ratio, resulting in an artificially high value. In light of this, to reflect the actual value, we purely calculate Hit ratio as the percentage of successful modifications without excluding modified molecules that are chemically invalid, among number of given molecules, as mathematically formulated in (3).
Appendix A2 Analysis
We conduct an analysis to understand the impact of the maximum number of iterations N in the iterative paradigm by growing N to larger limits. Notably, the procedure can be early stopped if the modified molecule meets the objective, as precisely depicted in Algorithm 1, growing N means growing the number of trials. Table 4 reveals that $\text{Re}^{3}$ is beneficial from growing N, gains improved performance as N gets larger and consistently outperforms the previous ReDF across both single- and multi-property objectives with 2 thresholds. flushbottom